OpenStudy (imtiaz7):

A train starts from rest. It moves through 1km in 100s with uniform acceleration. What will be its speed at the end of 100s. @mary.rojas @xxAshxx

OpenStudy (mary.rojas):

It starts from rest. So, u = 0 Distance traveled = s = 1 km= 1000 m Time of travel = t = 100 sec Substituting in s = u*t + (1/2)*a*t^2 (where a is the acceleration): 1000 = 0*t + (1/2)*a*100^2 1000 = a*5000 a = 0.2 m/sec^2 Next, substituting in v^2 - u^2 = 2*a*s (where v is the speed at the end of 100 sec): v^2 - 0^2 = 2*0.2*1000 => v^2 = 400 => v = 20 m/sec So, speed at the end of 100 sec is 20 m/sec. FROM http://answers.yahoo.com/question/index?qid=20121225234317AAGQbqp

OpenStudy (anonymous):

@imtiaz7 sorry,

OpenStudy (imtiaz7):

@xxAshxx no problem nothing anyone in world who is master in all subjects.

OpenStudy (anonymous):

@imtiaz7 lol...:-)